# Palmprint De-Identification Using Diffusion Model for High-Quality and Diverse Synthesis

### v1

@@ Abstract
Palmprint recognition techniques have advanced significantly in recent years, enabling reliable recognition even when palmprints are captured in uncontrolled or challenging environments. However, this strength also introduces new risks, as publicly available palmprint images can be misused by adversaries for malicious activities. Despite this growing concern, research on methods to obscure or anonymize palmprints remains largely unexplored. Thus, it is essential to develop a palmprint de-identification technique capable of removing identity-revealing features while retaining the image’s utility and preserving non-sensitive information. In this paper, we propose a training-free framework that utilizes pre-trained diffusion models to generate diverse, high-quality palmprint images that conceal identity features for de-identification purposes. To ensure greater stability and controllability in the synthesis process, we incorporate a semantic-guided embedding fusion alongside a prior interpolation mechanism. We further propose the de-identification ratio, a novel metric for intuitive de-identification assessment. Extensive experiments across multiple palmprint datasets and recognition methods demonstrate that our method effectively conceals identity-related traits with significant diversity across de-identified samples. The de-identified samples preserve high visual fidelity and maintain excellent usability, achieving a balance between de-identification and retaining non-identity information.
Index Terms:
Palmprint recognition, De-identification, Palmprint synthesis, Privacy protection, Diffusion models.
I Introduction
@@
An overview of the palmprint de-identification and recognition pipeline is illustrated in Fig. 1. As shown on the left side, the goal of the de-identified image is 1) to increase the feature distance from the original in the identity space, thereby denying recognition; 2) preserving the image’s quality and utility as closely as possible to the original.
@@
Extensive experiments conducted across several palmprint datasets and recognition systems, including both controlled and challenging scenarios, demonstrate that the proposed method offers strong de-identification capabilities. It effectively preserves non-identity-related information and usability. Additionally, our method exhibits high output diversity, producing distinctly different de-identified versions from the same palmprint input. Comparative analyses against conventional anonymization techniques, such as masking, blurring, and pixelating, highlight the superiority of our approach in striking a balance between identity concealment and image utility.
@@
4. Experimental results on diverse palmprint datasets and recognition methods demonstrate the effectiveness, generality, and robustness of our approach. The high diversity of generated de-identified outputs mitigates the risk of inverse inference.
The rest of the paper is organized as follows: Section II reviews related work, Section III details the diffusion model and proposed framework, Section IV presents experimental evaluations, and Section V concludes with a summary and future research directions.
II Related Works
@@ II-A Palmprint Recognition
II-B Biometrics De-identification
Biometric de-identification, also known as anonymization, refers to the process of obscuring identity-specific information in biometric data while preserving non-identifying attributes to maintain usability [26]. This typically involves modifying or replacing personal identifiers in a way that conceals sensitive information from public access. The de-identification pipeline, illustrated in Fig. 1, using palmprint as an example, aims to increase the disparity between the original and de-identified features to prevent successful recognition while simultaneously preserving the visual quality and utility of the de-identified sample.
@@ II-B Biometrics De-identification
III Methodology
III-A Preliminary
III-A1 Stable Diffusion
@@
III-A2 Paint by Example
Paint by Example [20] is an exemplar-based image inpainting diffusion model built upon the Stable Diffusion (SD) framework. It is trained in a self-supervised manner to reconstruct masked regions of an image by leveraging both a background image and an exemplar image from a same image as conditioning signals. Specifically, the model requires a masked background image and an exemplar image, using the contextual cues from the background and the semantic-guided embedding from the exemplar to fill in the missing region. Importantly, the exemplar image undergoes a semantic compression process through a bottleneck architecture composed of CLIP and a multilayer perceptron (MLP), ensuring that only high-level semantic features, not detailed identity-specific information, are retained. This prevents direct leakage of exemplar details into the inpainting result.
@@ III-A2 Paint by Example
III-B Palmprint De-identification
III-B1 Overall Pipeline of Our Method
@@
In the preparation step, all required components are extracted from a single hand image using a keypoint detection model and a segmentation model (Section III-B2). Two regions of interest, a small ROI (sROI) and a medium ROI (mROI), are isolated to serve as exemplar inputs to extract semantic-guided embedding (SGE). Simultaneously, a masked background image is obtained for constraining the inpainting process. Both the original palmprint and the masked background are essential inputs for the subsequent de-identification step.
In the de-identification step, key components, including CLIP and MLP, the VAE Encoder, and the diffusion model, are inherited from the Paint by Example framework. The CLIP&MLP module functions as a semantic bottleneck, extracting highly compressed SGE from the sROI and mROI. Both SGEs are then fused into a unified representation via a semantic-guided embedding fusion strategy (Section III-B3), capturing richer and complementary identity-independent features.
Meanwhile, the VAE encoder transforms both the original and background images into their respective latent maps. A prior interpolation strategy is then employed to compute a new latent map that lies between the original and background latent maps. This interpolated latent map acts as a controllable modulator, enabling a balance between effective de-identification and the preservation of image quality and usability (refer to Section III-B4).
Finally, the diffusion model synthesizes the masked region by integrating the fusion SGE and conditioning on the interpolated latent (refer to Section III-B5). The result is a high-quality, realistic palmprint image with de-identified textures that maintain usability while obscuring identity traits. Each component of this pipeline is elaborated in the subsequent sections.
III-B2 Keypoint Detection and Segmentation
@@
III-B3 Semantic-guided Embedding Fusion
In the original Paint by Example framework, only a single image is used as the exemplar. However, this setup is not tailored for palmprint generation, often leading to unpredictable or suboptimal results. To address this limitation, we introduce a simple yet effective fusion strategy that integrates SGE from exemplar ROIs at multiple scales. This multi-scale fusion enhances both the reliability and stability of the guidance signal. The fusion process is formalized as follows:
@@ III-B3 Semantic-guided Embedding Fusion
Since SGE functions as a semantic representation, our fusion strategy effectively seeks a balanced midpoint among these representations. By fusing SGE across multiple scales, this approach mitigates the influence of noise or distortions that may arise from any single exemplar. As a result, the fusion process enhances the stability and quality of the generated outputs, producing more consistent and desirable results.
III-B4 Prior Interpolation
In addition to exemplar semantic guidance, the inpainting results are also influenced by the latent representation of the background. When the background contains undesirable context, it can adversely affect the final output. A natural solution is to incorporate prior knowledge from the original image to guide the generation. However, this must be done carefully; excessive prior information risks compromising the de-identification objective. To address this trade-off, we propose a prior interpolation strategy that allows fine-grained control over the balance between visual quality and identity obfuscation. The prior interpolation mechanism is defined as follows:
@@ III-B4 Prior Interpolation
III-B5 De-identified Palmprint Synthesis
To synthesize de-identified palmprints, the fused SGE is injected into each block of the U-Net within the diffusion model using a cross-attention mechanism [33], progressively steering the denoising and hence generation process. Simultaneously, the interpolated latent representation and its corresponding mask are fed into the U-Net, along with random noise. During generation, the unmasked regions of the latent remain unchanged, preserving their original structure, while the masked regions are regenerated from a random noise. This regeneration is both guided by the fused SGE and constrained by the context of the surrounding unmasked areas, ensuring semantic consistency and structural realism.
III-C Evaluation Metrics
@@
where $\mu_{n}$, and $\sigma_{n}$ denote the mean and the standard deviation of distribution $n$, respectively.
where $n\in\{1,2\}$.
The DIR is designed to evaluate the distribution of de-identified match scores by referencing both genuine and imposter matching distributions. This enables a comprehensive, multidimensional assessment of de-identification performance. Formally, the DIR can be defined as follows:
@@
Generally, the closer the DIR is to 100%, the more effective the de-identification. However, values exceeding 100% should be avoided, as they indicate that the de-identified features deviate even more than typical imposters, possibly falling outside the expected distribution of normal samples and introducing recognition instability.
@@ IV Experiments
Datasets: Five contactless hand datasets are included in our experiments, as detailed in Tab. I, and their visual presentation is shown in Fig. 4.
TABLE I: Datasets introduction. The details of five contactless hand datasets.
@@
IV-A De-identification Performance
We selected two distinct palmprint recognition approaches for evaluation: MTCC, a hand-crafted feature-based method [4], and CCNet, a deep learning-based model [5]. Although our primary focus lies on palmprint regions, we also extended our analysis to include full-hand recognition by testing EEHNet [8]. Tab. II presents the comparative performance across various datasets and recognition methods. In the table, the symbols ↑ and ↓ indicate whether a higher or lower value is preferable, respectively.
TABLE II: De-identification performance (%). The symbols ↑ (↓) indicate that the higher (lower) value is better.
@@ IV-A De-identification Performance
As summarized in Tab. II, our de-identification method demonstrates strong performance across various recognition models and datasets. Specifically, in the hand-crafted MTCC method, our approach achieves a minimum of 85.63% DIR and a maximum accuracy of 10.91%. In contrast, for CCNet, the DIR reaches 68.41%, while accuracy peaks at 27.17%. Notably, on the PolyU dataset, we observe a favorable DIR of 71.04% and a peak accuracy of just 0.23%, despite a significantly low RR of 8.76%. This discrepancy highlights that DIR offers a more reliable measure of de-identification efficacy compared to RR.
@@
IV-B Quality of De-Identified Palmprints
@@
To quantitatively assess the quality of the de-identified images, we evaluated their similarity to the original images using five established metrics: SSIM, MS-SSIM, PSNR (where higher values indicate better quality), LPIPS and FID (where lower values are preferred). The results, presented in Tab. III, confirm that our method introduces minimal distortion while preserving the original style and structural consistency. This balance between visual realism and subtle identity modification underscores the strength of our de-identification strategy.
TABLE III: De-identified image quality. The symbols ↑ (↓) indicate that the higher (lower) value is better.
@@
IV-C Usability after De-identification
Usability plays a critical role in evaluating de-identification methods. An ideal approach should effectively remove identity-related features while preserving the image’s utility for other downstream tasks. In our study, we assess five key usability aspects: hand detection, 2D and 3D keypoint detection provided by Wilor [31], as well as ROI localization and hand segmentation provided by PKLNet [9].
Usability is quantified by computing the mean absolute error between the outputs generated from the original and de-identified images. To facilitate consistent comparison, all differences are normalized to a [0–100]scale, where lower values indicate better preservation of usability.
As shown in Tab. IV, our method maintains a usability difference of less than 1% in most scenarios, indicating that the de-identified images remain highly suitable for various non-identification tasks. This demonstrates that our framework achieves a desirable balance between identity removal and functional integrity.
TABLE IV: Usability differences of de-identification ↓ (%). The range of the value is [0-100]%, and the lower value is the better.
Dataset Hand Detection 2D Keypoint Detection 3D Keypoint Detection ROI Localization Hand Segmentation
@@ D. Comparison with Traditional Anonymization Methods:
Quantitative results for all experiments are presented in Tab. V, Tab. VI, and Tab. VII, which report performance in terms of de-identification accuracy, image quality, and usability. Horizontal lines delineate each experimental setting, and the best result within each segment is highlighted in bold. Visual outcomes are provided in Fig. 7, clearly illustrating the qualitative differences across configurations.
TABLE V: The de-identification performance of ablation study (%). The symbols ↑ (↓) indicate that the higher (lower) value is better, and the bold means the best result in the corresponding segment.
@@
s 97.54 97.00 6.65 59.42 69.54 0.42
m 97.77 97.08 6.58 57.73 69.94 0.31
f 98.00 97.68 6.38 61.23 70.35 0.23
s+m 97.50 96.85 7.15 56.00 69.35 0.50
s+f 97.73 97.05 6.62 57.12 69.58 0.35
m+f 97.88 97.26 6.54 58.31 70.19 0.31
s+m+f 97.58 96.99 6.81 56.38 69.59 0.31
$\alpha$=0.1 95.42 94.79 10.35 50.46 68.41 0.27
$\alpha$=0.2 91.46 92.50 15.46 44.46 67.29 0.23
$\alpha$=0.3 85.00 89.87 22.96 38.19 65.85 0.23
Masking 100.00 125.31 0.31 99.46 74.36 0.12
Blurring 0.04 34.94 100.00 6.27 46.89 0.23
Pixelating 3.38 44.83 97.92 3.04 47.18 0.12
TABLE VI: Image quality of ablation study. The symbols ↑ (↓) indicate that the higher (lower) value is better, and the bold means the best result in the corresponding segment.
@@
TABLE VII: Usability differences of ablation study ↓ (%). The range of the value is [0-100]%, and the lower value is the better. The bold means the best result in the corresponding segment.
Dataset Hand Detection 2D Keypoint Detection 3D Keypoint Detection ROI Localization Hand Segmentation
@@
Part C focused on tuning the prior interpolation factor $\alpha$ in the latent conditioning map under the s+m fusion strategy. When $\alpha$ was set to 0.1, the framework achieved its strongest de-identification capability, effectively suppressing identifiable traits. As $\alpha$ increased, more prior knowledge was incorporated into the generation process, which improved image quality and usability but weakened anonymization strength. Based on these observations, $\alpha$ = 0.1 was identified as the most balanced setting, offering robust de-identification while avoiding visually implausible artifacts. This value was adopted as the default interpolation factor in our experiments.
@@
V Conclusions and Future Works

### v2

# Palmprint De-Identification via Diffusion Model for High-Quality and Diverse Synthesis

@@ Abstract
Palmprint recognition techniques have advanced significantly in recent years, enabling reliable recognition even when palmprints are captured in uncontrolled or challenging environments. However, this strength also introduces new risks, as publicly available palmprint images can be misused by adversaries for malicious activities. Despite this growing concern, research on methods to obscure or anonymize palmprints remains largely unexplored. Thus, it is essential to develop a palmprint de-identification technique capable of removing identity-revealing features while retaining the image’s usability and preserving non-sensitive information. In this paper, we propose a training-free framework that utilizes pre-trained diffusion models to generate diverse, high-quality palmprint images that conceal identity features for de-identification purposes. To ensure greater stability and controllability in the synthesis process, we incorporate a semantic-guided embedding fusion alongside a prior interpolation mechanism. We further propose the de-identification ratio, a novel metric for intuitive de-identification assessment. Extensive experiments across multiple palmprint datasets and recognition methods demonstrate that our method effectively conceals identity-related traits with significant diversity across de-identified samples. The de-identified samples preserve high visual fidelity and maintain excellent usability, achieving a balance between de-identification and retaining non-identity information.
keywords:
Palmprint recognition , De-identification , Palmprint synthesis , Privacy protection , Diffusion models
journal: Pattern Recognition
[…]
[1]organization=PAMI Research Group, Department of Computer and Information Science, University of Macau,addressline=Taipa, city=Macau, postcode=999078, country=China
[…]
[2]organization=School of Electrical and Electronic Engineering, College of Engineering, Yonsei University,addressline=50 Yonsei-ro, city=Seodaemun-gu, postcode=120749, state=Seoul, country=Republic of Korea
[…]
[3]organization=School of Software, Nanchang Hangkong University,addressline=696 Fenghe Nan Avenue, city=Nanchang, postcode=330063, state=Jiangxi, country=China
[…]
[4]organization=College of Computer Science, Sichuan University,city=Chengdu, postcode=610045, state=Sichuan, country=China
1 Introduction
@@
An overview of the palmprint de-identification and recognition pipeline is illustrated in Fig. 1. As shown on the left side, the goal of the de-identified image is 1) to increase the feature distance from the original in the identity space, thereby denying recognition; 2) preserving the image’s quality and usability as closely as possible to the original.
@@
Extensive experiments conducted across several palmprint datasets and recognition systems, including both controlled and challenging scenarios, demonstrate that the proposed method offers strong de-identification capabilities. It effectively preserves non-identity-related information and usability. Additionally, our method exhibits high output diversity, producing distinctly different de-identified versions from the same palmprint input. Comparative analyses against conventional anonymization techniques, such as masking, blurring, and pixelating, highlight the superiority of our approach in striking a balance between identity concealment and image usability.
@@
4. Experimental results on diverse palmprint datasets and recognition methods demonstrate the effectiveness, generality, and robustness of our approach. The generated outputs exhibit high diversity, which helps mitigate the risk of inverse inference.
The rest of the paper is organized as follows: Section 2 reviews related work, Section 3 details the diffusion model and proposed framework, Section 4 presents experimental evaluations, Section 5 discusses key design choices, benefits, and flexibility of the proposed framework, and Section 6 concludes with a summary and future research directions.
2 Related Works
@@ 2.1 Palmprint Recognition
2.2 Biometrics De-identification
Biometric de-identification, also known as anonymization, refers to the process of obscuring identity-specific information in biometric data while preserving non-identifying attributes to maintain usability [26]. This typically involves modifying or replacing personal identifiers in a way that conceals sensitive information from public access. The de-identification pipeline, illustrated in Fig. 1, using palmprint as an example, aims to increase the disparity between the original and de-identified features to prevent successful recognition while simultaneously preserving the visual quality and usability of the de-identified sample.
@@ 2.2 Biometrics De-identification
3 Methodology
3.1 Preliminary
3.1.1 Stable Diffusion
@@
3.1.2 Paint by Example
Paint by Example [20] is an exemplar-based image inpainting diffusion model built upon the Stable Diffusion (SD) framework. It is trained in a self-supervised manner to reconstruct masked regions of an image by leveraging both a background image and an exemplar image from a same image as conditioning signals. Specifically, the model takes a masked background image and an exemplar image as inputs, leveraging contextual cues from the background and semantic-guided embeddings from the exemplar to complete the missing region. In addition, the contextual cues are encoded into latent maps through the VAE encoder. Importantly, the exemplar image undergoes a semantic compression process through a bottleneck architecture composed of CLIP and a multilayer perceptron (MLP), ensuring that only high-level semantic features, not detailed identity-specific information, are retained. This prevents direct leakage of exemplar details into the inpainting result.
@@ 3.1.2 Paint by Example
3.2 Palmprint De-identification
3.2.1 Overall Pipeline of Our Method
@@
In the preparation step, all required components are extracted from a single hand image using a keypoint detection model and a segmentation model (Section 3.2.2). Two regions of interest, a small ROI (sROI) and a medium ROI (mROI), are isolated to serve as exemplar inputs to extract semantic-guided embedding (SGE). Simultaneously, a masked background image is obtained for constraining the inpainting process. Both the original palmprint and the masked background are essential inputs for the subsequent de-identification step.
In the de-identification step, key components, including CLIP and MLP, the VAE Encoder, and the diffusion model, are inherited from the Paint by Example framework. The CLIP&MLP module functions as a semantic bottleneck, extracting highly compressed SGE from the sROI and mROI. Both SGEs are then fused into a unified representation via a semantic-guided embedding fusion strategy (Section 3.2.3), capturing richer and complementary identity-independent features.
Meanwhile, the VAE encoder transforms both the original and background images into their respective latent maps. A prior interpolation strategy is then employed to compute a new latent map that lies between the original and background latent maps. This interpolated latent map acts as a controllable modulator, enabling a balance between effective de-identification and the preservation of image quality and usability (refer to Section 3.2.4).
Finally, the diffusion model synthesizes the masked region by integrating the fusion SGE and conditioning on the interpolated latent (refer to Section 3.2.5). The result is a high-quality, realistic palmprint image with de-identified textures that maintain usability while obscuring identity traits. Each component of this pipeline is elaborated in the subsequent sections.
3.2.2 Keypoint Detection and Segmentation
@@
3.2.3 Semantic-guided Embedding Fusion
In the original Paint by Example framework, only a single image is used as the exemplar. Semantic-guided embedding (SGE) compresses the exemplar image through a narrow bottleneck before conditioning the diffusion U-Net, thereby reducing identity leakage—such as fine-grained palmprint textures—while retaining higher-level semantic cues, such as overall palmprint appearance. Nevertheless, this mechanism is not specifically optimized for palmprint generation and may yield unpredictable or suboptimal results, often caused by inaccurate semantic representation or interference from semantic noise. To address this limitation, we introduce a simple yet effective SGE fusion strategy that integrates SGE from exemplar ROIs at multiple scales. This multi-scale fusion enhances both the reliability and stability of the guidance signal. The fusion process is formalized as follows:
@@ 3.2.3 Semantic-guided Embedding Fusion
Since SGE functions as a semantic representation, our fusion strategy effectively seeks a balanced midpoint among these representations. By fusing SGE across multiple scales, the proposed approach mitigates noise or distortions originating from a single exemplar and reinforces semantic consistency. As a result, the fusion process enhances the stability and quality of the generated outputs, producing more consistent and desirable results.
3.2.4 Prior Interpolation
In addition to exemplar semantic guidance, the inpainting results are also influenced by the latent representation of the background. When the background contains undesirable context, it can adversely affect the final output. A natural solution is to incorporate prior knowledge from the original image to guide the generation. By incorporating prior structural information from the original image, the generation process can be stabilized and constrained by the given structural context. However, this must be done carefully; excessive prior information risks compromising the de-identification objective. To address this trade-off, we propose a prior interpolation strategy that allows fine-grained control over the balance between visual quality and identity obfuscation. The prior interpolation mechanism is defined as follows:
@@ 3.2.4 Prior Interpolation
Here, larger $\alpha$ values favor usability and perceptual fidelity, whereas smaller $\alpha$ values enhance identity suppression at the potential cost of realism. Intuitively, minimizing identity leakage require a small $\alpha$ for prior context guidance. In our experiments, we set $\alpha=0.1$ as a balanced point, achieving a desirable trade-off between identity similarity, perceptual fidelity, and usability across datasets.
3.2.5 De-identified Palmprint Synthesis
The generation of de-identified palmprints in the diffusion model follows the same principle as Paint by Example [20]. To synthesize de-identified palmprints, the fused SGE is injected into each block of the U-Net within the diffusion model using a cross-attention mechanism [33], progressively steering the denoising and hence generation process. Simultaneously, the interpolated latent representation and its corresponding mask are fed into the U-Net, along with random noise. During generation, the unmasked regions of the latent remain unchanged, preserving their original structure, while the masked regions are regenerated from a random noise. This regeneration is both guided by the fused SGE and constrained by the context of the surrounding unmasked areas, ensuring semantic consistency and structural realism.
3.3 Evaluation Metrics
@@
where $\mu_{n}$ and $\sigma_{n}$ denote the mean and standard deviation of distribution $n$ $(n\in{1,2})$, respectively.
The DIR is designed to evaluate the distribution of de-identified match scores by referencing both genuine and imposter matching distributions, which serves as an index indicating whether the de-identified matches are closer to the genuine or imposter distribution. Ideally, de-identified matches should lie closer to the imposter distribution and farther from the genuine distribution. In particular, DIR evaluates de-identification performance from the matching distribution perspective, avoiding sensitivity to threshold selection or the influence of abnormal matches. This enables a comprehensive, multidimensional assessment of de-identification performance. Formally, the DIR can be defined as follows:
@@
For interpretability, we define qualitative bands of identity suppression effectiveness: DIR >> 80% denotes high suppression effectiveness, values in the range of 60–80% indicate moderate suppression effectiveness, and DIR << 60% reflects limited suppression effectiveness. Generally, the closer the DIR is to 100%, the more effective the de-identification. However, values exceeding 100% should be avoided, as they indicate that the de-identified features deviate even more than typical imposters, possibly falling outside the expected distribution of normal samples and introducing recognition instability.
@@ 4 Experiments
Datasets: Five contactless hand datasets are included in our experiments, as detailed in Tab. 1, and their visual presentation is shown in Fig. 4.
Table 1: Datasets introduction. The details of five contactless hand datasets.
@@
4.1 De-identification Performance
We selected two distinct palmprint recognition approaches for evaluation: MTCC, a hand-crafted feature-based method [4], and CCNet, a deep learning-based model [5]. Although our primary focus lies on palmprint regions, we also extended our analysis to include full-hand recognition by testing EEHNet [8]. Tab. 2 presents the comparative performance across various datasets and recognition methods. In the table, the symbols ↑ and ↓ indicate whether a higher or lower value is preferable, respectively.
Table 2: De-identification performance (%). The symbols ↑ (↓) indicate that the higher (lower) value is better.
@@ 4.1 De-identification Performance
As summarized in Tab. 2, our de-identification method demonstrates strong performance across various recognition models and datasets. Specifically, in the hand-crafted MTCC method, our approach achieves a minimum of 85.63% DIR and a maximum accuracy of 10.91%. In contrast, for CCNet, the DIR reaches 75.49% at least, while accuracy peaks at 27.17% at most. Notably, on the REST dataset, we observe a favorable DIR of 84.40% and a peak accuracy of just 9.19%, despite a low RR of 51.08%. This discrepancy highlights that DIR offers a more reliable measure of de-identification efficacy compared to RR.
@@
4.2 Quality of De-Identified Palmprints
@@
To quantitatively assess the quality of the de-identified images, we evaluated their similarity to the original images using five established metrics: SSIM, MS-SSIM, PSNR (where higher values indicate better quality), LPIPS and FID (where lower values are preferred). The results, presented in Tab. 3, confirm that our method introduces minimal distortion while preserving the original style and structural consistency. This balance between visual realism and subtle identity modification underscores the strength of our de-identification strategy.
Table 3: De-identified image quality. The symbols ↑ (↓) indicate that the higher (lower) value is better.
@@
4.3 Usability after De-identification
Usability plays a critical role in evaluating de-identification methods. An ideal approach should effectively remove identity-related features while preserving the image’s usability for other downstream tasks. In our study, we assess five key usability aspects: hand detection, 2D and 3D keypoint detection provided by Wilor [31], as well as ROI localization and hand segmentation provided by PKLNet [9].
Usability is quantified by computing the mean absolute error between the outputs generated from the original and de-identified images. To facilitate consistent comparison, all differences are normalized to a [0–100]% scale, where lower values indicate better preservation of usability.
As shown in Tab. 4, our method maintains a usability difference of less than 1% in most scenarios, indicating that the de-identified images remain highly suitable for various non-identification tasks. This demonstrates that our framework achieves a desirable balance between identity removal and functional integrity.
Table 4: Usability differences of de-identification ↓ (%). The range of the value is [0-100]%, and the lower value is the better.
Dataset
Hand
Detection
2D Keypoint
Detection
3D Keypoint
Detection
ROI
Localization
Hand
Segmentation
@@ D. Comparison with Traditional Anonymization Methods:
Quantitative results for all experiments are presented in Tab. 5, Tab. 6, and Tab. 7, which report performance in terms of de-identification accuracy, image quality, and usability. Horizontal lines delineate each experimental setting, and the best result within each segment is highlighted in bold. Visual outcomes are provided in Fig. 7, clearly illustrating the qualitative differences across configurations.
Table 5: The de-identification performance of ablation study (%). The symbols ↑ (↓) indicate that the higher (lower) value is better, and the bold means the best result in the corresponding segment.
@@
s 97.54 97.00 6.65 97.23 83.64 8.77
m 97.77 97.08 6.58 97.65 84.15 8.35
f 98.00 97.68 6.38 98.23 85.04 7.81
s+m 97.50 96.85 7.15 97.15 83.61 8.77
s+f 97.73 97.05 6.62 97.08 83.92 9.04
m+f 97.88 97.26 6.54 98.00 84.39 8.31
s+m+f 97.58 96.99 6.81 97.38 83.89 8.81
$\alpha$=0.1 95.42 94.79 10.35 96.42 82.55 10.12
$\alpha$=0.2 91.46 92.50 15.46 95.65 81.45 11.81
$\alpha$=0.3 85.00 89.87 22.96 94.15 80.07 15.00
Masking 100.00 125.31 0.31 100.00 82.79 0.38
Blurring 0.04 34.94 100.00 12.23 48.79 88.73
Pixelating 3.38 44.83 97.92 13.27 50.13 94.04
Table 6: Image quality of ablation study. The symbols ↑ (↓) indicate that the higher (lower) value is better, and the bold means the best result in the corresponding segment.
@@
Table 7: Usability differences of ablation study ↓ (%). The range of the value is [0-100]%, and the lower value is the better. The bold means the best result in the corresponding segment.
Dataset
Hand
Detection
2D Keypoint
Detection
3D Keypoint
Detection
ROI
Localization
Hand
Segmentation
@@
Part C focused on tuning the prior interpolation factor $\alpha$ in the latent conditioning map under the s+m fusion strategy. When $\alpha$ was set to 0.1, the framework achieved its strongest de-identification capability, effectively suppressing identifiable traits. As $\alpha$ increases, the interpolated latent moves closer to the original image, incorporating more prior knowledge into the generation process. As a result, the image quality and usability become more similar to the original image, but anonymization strength is correspondingly reduced. Based on these observations, $\alpha$ = 0.1 was identified as the most balanced setting, offering robust de-identification while avoiding visually implausible artifacts. This value was adopted as the default interpolation factor in our experiments.
@@
5 Discussion
We adopt a training-free, optimization-free design for three key reasons: privacy and governance (avoiding fine-tuning on biometric corpora minimizes the risk of model memorization and avoids consent and data-retention requirements), cross-domain generality (an off-the-shelf generator reduces overfitting to specific datasets or capture protocols), and reproducibility cost (results can be reproduced without large palmprint datasets or extensive computation). While this approach may slightly compromise modality-specific texture fidelity, it ensures stronger privacy and broader applicability. Importantly, the proposed plug-and-play mechanisms—SGE fusion and prior interpolation—provide flexibility, allowing the framework to be readily adjusted or reconfigured for different scenarios simply by tuning guidance or parameters, without any additional training or data collection. Overall, this paradigm achieves a practical balance between privacy, robustness, and accessibility, while leaving room for principled, privacy-aware adaptations in future work.
6 Conclusions and Future Works